### v1

# AutoCrawler : A Progressive Understanding Web Agent for Web Crawler Generation

Abstract
Web automation is a significant technique that accomplishes complicated web tasks by automating common web actions, enhancing operational efficiency, and reducing the need for manual intervention. Traditional methods, such as wrappers, suffer from limited adaptability and scalability when faced with a new website. On the other hand, generative agents empowered by large language models (LLMs) exhibit poor performance and reusability in open-world scenarios. In this work, we introduce a crawler generation task for vertical information web pages and the paradigm of combining LLMs with crawlers, which helps crawlers handle diverse and changing web environments more efficiently. We propose AutoCrawler , a two-stage framework that leverages the hierarchical structure of HTML for progressive understanding. Through top-down and step-back operations, AutoCrawler can learn from erroneous actions and continuously prune HTML for better action generation. We conduct comprehensive experiments with multiple LLMs and demonstrate the effectiveness of our framework. [^1] Resources of this paper can be found at https://github.com/EZ-hwh/AutoCrawler
22footnotetext: Corresponding authors.
1 Introduction
Web automation refers to the process of programmatically interacting with web-based applications or websites to execute tasks that would typically require human intervention. Web automation streamlines repetitive and time-consuming tasks, significantly enhancing efficiency, accuracy, and scalability across diverse online processes.
In traditional web automation, methods predominantly rely on wrappers, which are scripts or software specifically designed to extract data from predetermined websites or pages. This approach is characteristic of a closed-world scenario, where the automation system only interacts with a pre-defined, limited set of websites or pages and does not extend beyond this specified domain. Consequently, these traditional methods face limitations in adaptability and scalability, struggling to function effectively when encountering new or altered website structures. Given these limitations, both rule-based wrappers and auto wrappers Bronzi et al. (2013), despite their differences, share a common dependency on manually annotated examples for each website Gulhane et al. (2011). For example, more than 1,400 annotations of webpages are used for information extraction Lockard et al. (2019).
Figure 1: Top: HTML have a hierarchical structure DOM tree; Down: Existing web automation framework: Olive arrows refer to handcraft/LLMs prompting process, Violet arrows refer to parser.
The advent of LLMs has revolutionized web automation by introducing advanced capabilities such as planning, reasoning, reflection, and tool use. Leveraging these capabilities, web automation employs LLMs to construct generative agents that can autonomously navigate, interpret, and interact with web content. This effectively solves open-world web-based tasks through sophisticated language understanding and decision-making processes. However, despite these advancements, this paradigm faces two major issues. On one hand, existing web agent frameworks often demonstrate poor performance, with a success rate mentioned as 2.0 Deng et al. (2023) on open-world tasks. On the other hand, a significant weakness encountered in this approach is its insufficient reusability. This implies that these agents are overly dependent on LLMs even when dealing with similar tasks, leading to low efficiency when managing a large volume of repetitive and similar webpages.
In this work, we propose a crawler generation task for vertical information web pages. The goal of this task is to automatically generate a series of predefined rules or action sequences to automatically extract target information. This task calls for a paradigm that LLMs generate crawlers. Compared to traditional wrappers, this paradigm can be quickly adjusted according to different websites and task requirements. This flexibility enables crawlers to handle diverse and changing web environments more efficiently. Compared to the generative agent paradigm, it introduces intermediate rules to enhance reusability and reduce the dependency on LLMs when dealing with similar tasks, thereby improving efficiency when handling a large number of web tasks.
Although LLMs possess strong webpage information extraction ablities, there are still the following challenges in generating crawlers for LLMs: First, LLMs are primarily pre-trained on massive corpora of cleansed, high-quality pure text, lacking exposure to markup languages such as HTML. As a result, LLMs exhibit a limited understanding of the complex structures and semantics inherent in HTML. Second, HTML, as a semi-structured data, contains elements of both structured (tags and attributes) and unstructured (textual content), concurrently encompassing multilayered information nested. It augments the complexity of crawler generation. Third, although LLMs excel in comprehending textual content, they still fall short in understanding the structural information of lengthy documents. This indicates a potential challenge in accurately capturing and utilizing the hierarchical structure inherent in long HTML documents.
Therefore, we introduce AutoCrawler , a two-stage framework designed to address the crawler generation task. An overview of AutoCrawler is presented in Figure 2. Our framework leverages the hierarchical structure of HTML for progressive understanding. Specifically, we propose a heuristic algorithm consisting of top-down and step-back operation based on LLMs. It first tries to refine down to the specific node in the DOM tree containing the target information, and then moves up the DOM tree when execution fails. This process can correct erroneous executions and progressively prune irrelevant parts of the HTML content until it successfully executes.
Our contributions can be summarized as follows:
  • We propose the web crawler generation task and the paradigm of utilizing LLMs for generating crawlers, and we make an analysis on extraction efficiency.
  • We introduce AutoCrawler , a two-phase framework with progressive understanding to generate executable action sequences.
  • Comprehensive experimental results demonstrate the effectiveness of our framework in the web crawler generation task.
2 Preliminaries
Figure 2: Our framework for the generation of crawler. Left: Progressive generation process, consists of a cycle of top-down and step-back to progressively generate an executable action sequence; Right: Synthesis process, generates a stable action sequence generated based on ones from seed websites.
In this section, we first define the crawler generation task and then present the dataset collection process and its corresponding evaluation metrics.
2.1 Task Formulation
First, we formulate our crawler generation task. Given a set of webpages on the same website $w\in\mathcal{W}$ describing a subject entity $s$ (also called a topic entity in the previous literature), and its corresponding predefined target attribute $r\in\mathcal{R}$, the task objective is to generate an executable rule/action sequence $\mathcal{A}$ to extract target information $o$ from all webpages.
2.2 Datasets
We adopt the semi-structure information extraction task as a testbed for the crawler generation task.
Swde
Hao et al. (2011) is a Structured Web Data Extraction dataset that contains webpages and golden labels from 80 websites in 8 domains, with 124,291 webpages. Each of the websites from the same domains focuses on 3-5 attributes in the webpages.
Extended Swde
Lockard et al. (2019) involves fine-grained manual annotation of 21 sites in 3 domains from Swde. While Swde contains an average of 4,480 triples for 3 predicates per website, the Extended Swde dataset averages 41K triples for 36 predicates per site.
Ds1
Omari et al. (2017) contains 166 annotated webpages from 30 real-life large-scale websites categorized into books, shopping, hotels, and movies.
We transform the dataset with the following settings. First, we design instructions for each of the domains, and for each of the attributes as the input information for LLMs [^2] Further details about the prompt is in Appendix D.1. Second, for each website in each domain, we sample 100 webpages as the whole test set. We consider the set of webpages on the same websites and the corresponding extraction instruction as a case. For example, for the ESPN websites [^3] https://global.espn.com/nba/ in NBA player domains, the sampled 100 detail webpage of players and the instruction Please extract the team of the player he plays now is a complete test case of our crawler generation task. Third, we pre-process the websites by removing those elements in a webpage that do not contribute to the semantics. We filter out all DOM element nodes with <script> and <style>, as well as delete all attributes in the element node except @class. We replace the original escape characters in the annotations to ensure consistency with the corresponding information on the web.
Ultimately, we collect three datasets containing 320 / 294 / 83 test cases, covering 32 / 221 / 11 different extraction tasks, and comprising a total of 61,566 webpages from 10 different domains.
2.3 Evaluation Metrics
A single generation from an LLM is capable of directly extracting value from web pages and generating sequences of actions. However, the existing evaluation schemes for web page extraction tasks still follow the traditional metrics of text information extraction tasks, namely precision, recall, and F1 score. They limit the assessment of methods for the crawler generation task to two aspects. First, it focuses on extraction with a single webpage, rather than considering the generalizability from the perspective of a collection of webpages. Second, it does not effectively measure the transferability when adopting the action sequence to other webpages.
To address this issue, we transform the traditional IE task evaluation into an executable evaluation. Based on the traditional IE evaluation on a collection of webpages, we categorize the executability of action sequences into the following six situations. Specifically, for each extraction task on a website, the result is classified based on the extraction result on precision, recall, and f1-score.
(1) Correct, both precision, recall and f1-score equal 1, which indicates the action sequence is precisely; (2) Precision(Prec.), only precision equals 1, which indicates perfect accuracy in the instances extracted following the action sequence, but misses relevant instances; (3) Recall(Reca.), only recall equals 1, which means that it successfully identifies all relevant instances in the webpage but incorrectly identifies some irrelevant instances; (4) Un-executable(Unex.), recall equals 0, which indicates that the action sequence fails to identify relevant instances; (5) Over-estimate(Over.), precision equals 0, which indicates that the action sequence extracts the instances while ground truth is empty; (6) Else: the rest of the situation, including partially extracting the information, etc.
Since the above classifications are mutually exclusive, we use the ratio metric to calculate the proportion of each result in our task.
$M_{R}=\frac{\#\ \text{case of situation}}{\#\ \text{total case}}$ (1)
We are more concerned with success rate, so for the Correct metric, higher values indicate a better proportion of generated execution paths; whereas for the Un-executable metric, lower values are preferable.
3 AutoCrawler
In this section, we describe our framework AutoCrawler for generating a crawler to extract specific information from semi-structured HTML. Our approach is divided into two phases: first, we adopt a progressive generation framework that utilizes the hierarchical structure of web pages; second, we employ a synthesis framework based on results from multiple web pages. The overall framework is presented in Figure 2.
1
Data: origin HTML code $h_{0}$, task instruction $I$, max retry times $d_{max}$
2
Result: Executable action sequence $\mathcal{A}_{seq}$ to extract the value in the HTML
3
4Initial history $\mathcal{A}_{seq}\leftarrow[],k=0$;
5while True do
6       if $k>d_{max}$ then break;
7
      // Top-down
8       $value,xpath\leftarrow\textbf{LLM}_{g}(h_{k},I)$;
9       $result\leftarrow\textbf{Parser}_{text}(h_{k},xpath)$;
10       if result == value then break;
11
      // Step-back
12       repeat
13             $xpath\leftarrow xpath+"/.."$;
14             $h_{k+1}\leftarrow\textbf{Parser}_{node}(h_{k},xpath)$;
15
16      until h contains value;
17      $\text{Append}(\mathcal{A}_{seq},xpath)$;
18       $k\leftarrow k+1$;
19
20 end while
21
return $\mathcal{A}_{seq}$
Algorithm 1 Algorithm for progressive understanding
3.1 Modeling
Unlike the wrapper method that generates an XPath, we model the crawler generation task as an action sequence generation task. In specific, we generate an action sequence $\mathcal{A}_{seq}$ that consists of a sequence of XPath [^4] https://en.wikipedia.org/wiki/XPath expression from a set of seed webpages (i.e., a small portion of webpages in the test case for generating the sequence).
$\mathcal{A}_{seq}=[\text{XPath}_{1},\text{XPath}_{2},...,\text{XPath}_{n}]$ (2)
where $n$ denotes the length of the action sequence. We execute the XPath in the sequence using the parser in order. In the sequence, all XPath expressions except the last one are used for pruning the webpage, and the last one is used for extracting the corresponding element value from the pruned web page.
Models Method Executable Evaluation IE Evaluation
Correct($\uparrow$) Prec Reca Unex.($\downarrow$) Over. Else Prec Reca F1
Closed-source LLMs
GPT-3.5-Turbo COT 36.75 8.83 6.71 43.46 0.71 3.53 89.45 50.43 47.99
Reflexion 46.29 11.66 2.83 37.10 0.71 1.41 94.67 55.85 55.10
AutoCrawler 54.84 11.83 8.96 19.35 1.08 3.94 85.85 73.34 69.20
[…]
Reflexion 33.12 6.56 4.06 52.50 0.63 3.12 87.45 42.75 40.88
AutoCrawler 42.81 11.87 4.69 34.38 1.25 5.00 85.70 57.54 54.91
[…]
Reflexion 67.50 13.75 4.37 10.94 0.94 2.50 93.28 82.76 82.40
AutoCrawler 71.56 14.06 5.31 4.06 0.63 4.37 92.49 89.13 88.69
Open-source LLMs
Mistral 7B COT 3.44 0.31 0.63 95.31 0.00 0.63 94.23 4.55 4.24
Reflexion 2.19 0.00 0.31 97.19 0.00 0.31 95.60 2.78 2.49
AutoCrawler 2.87 0.00 0.00 96.77 0.36 0.00 98.57 3.23 2.87
[…]
Reflexion 18.08 4.80 2.95 73.06 0.00 1.11 78.96 23.26 22.44
AutoCrawler 23.99 8.12 1.48 64.94 0.00 1.48 78.59 28.70 28.41
[…]
Reflexion 36.25 6.88 3.12 51.25 0.00 2.50 89.35 44.57 43.60
AutoCrawler 46.88 10.62 7.19 30.31 0.63 4.37 87.32 62.71 59.75
[…]
Reflexion 37.19 11.25 4.06 44.69 1.25 1.56 86.41 48.28 47.08
AutoCrawler 38.75 11.25 5.31 39.69 0.63 4.37 84.91 52.11 49.68
Table 1: The executable evaluation and IE evaluation of LLMs with three frameworks in Swde dataset. We examine 7 LLMs, including 3 closed-source LLMs and 4 open-source LLMs.
3.2 Progressive Generation
Dealing with the lengthy content and hierarchical structure of webpages, generating a complete and executable crawler in one turn is difficult. However, the HTML content is organized in a DOM tree structure, which makes it possible to prune irrelevant page components and hence, limit the length and height of the DOM tree to improve the performance of LLM generation.
Specifically, we perform a traversal strategy consisting of top-down and step-back operations. Top-down refers to starting from the root node of the current DOM tree, progressively refining down to the specific node containing the target information. Step-back refers to reassessing and adjusting selection criteria by moving up the DOM tree to choose a more reliable and broadly applicable node as a foundation for more consistent and accurate XPath targeting. At each step, we first employ a top-down operation, guiding the LLMs to directly write out the XPath leading to the node containing the target information and to judge whether the value extracted with XPath is consistent with the value it recognizes. If execution fails, then adopt a step-back operation to retreat from the failed node, ensuring the web page includes the target information, which is driven by LLMs. The detail is shown in Algorithm 1.
3.3 Synthesis
Although we gain an executable action sequence within the progressive generation process, there are still differences in the specific location of the target information and the structure between different web pages. The action sequence may collect XPath with specific characteristics in a single HTML and lose generalizability. To enhance the reusability of the action sequence, we propose a synthesis phase.
Specifically, we randomly select $n_{s}$ webpages from the test case as seed webpages. Then, we generate an action sequence for each of them. Subsequently, we execute multiple different action sequences to extract information from the seed webpages, respectively. We collect all action sequence and their corresponding results and then choose one action sequence that can extract all the target information in the webpages as the final action sequence.
4 Experiment
Intending to put AutoCrawler to practical use, we investigate the following research questions: 1) How is the overall performance of AutoCrawler in comparison to the current state-of-the-art crawler generation task? 2) How does our progressive understanding generation framework improve performance? What is its relationship to the size of LLM? 3) Can we rely entirely on LLMs for web automation? 4) In which scenarios do current frameworks still not perform well?
4.1 Experimental Settings & Evaluation Metrics
We conduct our experiment on various LLMs including closed-source LLMs: GPT-3.5-Turbo OpenAI (2022), Gemini ProTeam et al. (2023) and GPT4 OpenAI (2023) as well as open-source LLMs: Mistral-7B-Instruct-v0.2 Jiang et al. (2023), CodeLlama-34B-Instruct Rozière et al. (2024), Mixtral 8$\times$7B-Instruct-v0.1 Jiang et al. (2024) and Deepseek-Coder-33B-Instruct Guo et al. (2024). Furthermore, we apply different LLM-prompt-based web agents as our baselines, including ZS_COT Wei et al. (2023) and Reflexion Shinn et al. (2023) and AutoCrawler to them. Due to the limited-length context of LLMs, all experiments are conducted under zero-shot settings. The full prompts can be found in Appendix D.2.
We test them on three datasets: Swde Hao et al. (2011), Extend Swde Lockard et al. (2019) and Ds1 Omari et al. (2017). The experimental result of the last two can be found in Appendix A.1 and  A.2. We set the size of seed webpages $n_{s}=3$ and max retry times $d_{max}=5$.
In addition to the execution evaluation metrics described in Section 2.3, we also employ traditional evaluation metrics to more comprehensively assess the quality of different action sequences. Specifically, we adopt precision (P.), recall (R.), and macro-f1 (F1), which are calculated as the mean of the corresponding metrics for each case.
4.2 Main Results on Swde
Results in Table 3.1 show that: 1) With AutoCrawler generating action sequence, LLMs can achieve better performance. Compared to the COT and Reflexion baseline, our method performs a higher ratio of correct and lower ratio of un-executable. Also, it should be noted that Mixtral 8$\times$7B + AutoCrawler can outperform ChatGPT + Reflexion, indicating the superiority of AutoCrawler in the generation of executable action sequences in the crawler generation task. 2) Models with small parameter sizes have significant difficulties in understanding and writing executable paths, so they can be considered challenging to apply in this task. On the contrary, large-scale models demonstrate a more stable ability in instruction alignment, web structure comprehension, and reflection on execution results; 3) Traditional IE evaluation metrics cannot well describe the success rate of our task. Especially for the precision metric, it fails to reveal the performance gap among different methods with different models. This is because the extraction metrics only evaluate the results that have been extracted, ignoring that unexecutable or empty extractions also greatly damage the executability.
4.3 Generate with Golden Label
Models Method Executable Evaluation
Correct($\uparrow$) Prec Reca Unex.($\downarrow$) Over. Else
Closed-source LLMs
GPT-3.5-Turbo COT 41.70 12.92 7.38 35.42 0.74 1.85
Reflexion 47.23 16.24 2.21 33.21 0.37 0.74
AutoCrawler 56.89 19.43 5.65 13.43 0.71 3.89
[…]
Reflexion 35.31 9.38 6.88 43.75 1.56 3.12
AutoCrawler 45.31 13.44 6.25 30.31 1.25 3.44
[…]
Reflexion 71.25 7.19 4.69 14.37 0.94 1.56
AutoCrawler 75.31 10.94 4.37 4.06 0.63 4.69
Open-source LLMs
Mistral 7B COT 2.19 0.00 0.31 97.19 0.00 0.31
Reflexion 2.19 0.00 0.00 97.50 0.31 0.00
AutoCrawler 2.19 0.00 0.00 97.19 0.31 0.31
[…]
Reflexion 22.21 4.93 3.94 66.95 0.49 1.48
AutoCrawler 26.20 12.55 5.54 53.51 0.00 2.21
[…]
Reflexion 34.69 8.13 5.31 49.06 0.63 2.19
AutoCrawler 45.62 11.56 5.94 32.50 1.25 3.12
[…]
Reflexion 38.75 11.87 2.81 42.19 0.63 3.75
AutoCrawler 38.44 20.94 4.06 31.56 0.94 6.56
Table 2: The executable and IE evaluation with 7 LLMs on Swde dataset with golden label.
To better illustrate the effectiveness of our framework in generating executable action sequences, we compare the performance of the COT, Reflexion, and our framework, while giving the golden label of the instruction. By offering the same extraction targets, we can effectively detect the actual effects of different frameworks on generating action sequences.
Table 4.3 shows experimental results, from which we can have the following observations: 1) Our proposed progressive understanding framework still effectively enhances the model’s performance under this setting; 2) LLMs still suffer in accurately understanding web page contents with semi-structured markup languages, which illustrate the performance gap between Table 3.1 and Table 4.3; 3) Compared to closed-source LLMs, even provided with golden labels, Open-source LLMs are unable to achieve sustained performance improvement. This phenomenon demonstrates that the bottleneck for these models lies not in understanding the webpage content but in understanding the webpage’s hierarchical structure itself.
Models 1 2 3 4 5 Avg.
GPT4 214 61 13 18 10 1.57
GPT-3.5-Turbo 115 65 22 30 43 2.35
Gemini Pro 94 52 33 27 105 2.99
Mixtral 8$\times$7B 89 53 43 24 104 3.00
Mistral 7B 28 7 11 7 84 3.82
Deepseek-coder 137 70 55 29 23 2.14
CodeLlama 75 35 32 18 80 2.97
Table 3: Length of action sequence of AutoCrawler based on different LLMs in Swde dataset.
4.4 Further Study with AutoCrawler
The length of the action sequence is dependent on the LLMs capability.
To comprehensively explore the performance of different LLMs in understanding web page structure, we explore the impact of models on the number distribution of the steps. In particular, we collect all the action sequences and calculate the average steps of AutoCrawler with different LLMs. The experimental result is reported in Table 3.
We observe that AutoCrawler with stronger LLMs generate fewer lengths of action sequence. AutoCrawler with GPT4 generates 1.57 steps on average, while the AutoCrawler with Mistral 7B generates 3.82 steps on average. This phenomenon can be interpreted as more powerful models having a better understanding of the web page hierarchical structure, thus being able to accurately output the appropriate XPaths in longer/deeper web pages, thereby reducing the number of steps.
Figure 3: The curve on length and height compression ratio in Swde dataset.
The "U" curve of compression ratio
We define the length of HTML as the number of tokens in the HTML, and its height as the height of the DOM tree represented by the HTML. we define the compression ratio of length and height as the ratio of the length/height of the original web page to that of the web page after being pruned by AutoCrawler .
$Compression_{L}=\frac{\#\text{tokens of new HTML}}{\#\text{tokens of origin HTML}}$ (3)
$Compression_{H}=\frac{\#\text{height of new HTML}}{\#\text{height of origin HTML}}$
We calculate their compression ratio of the Correct case and rank LLMs based on their performance. Figure 3 shows the result. It is interesting to note that there is a "U" curve on both the length and height compression ratios. This phenomenon can be explained from two aspects: on one hand, when LLM is powerful, it can generate the correct XPath without the process of step-back to re-accessing the sub-DOM tree; on the other hand, when the model is weak, it is unable to effectively understand the hierarchical structure of web page, and thus cannot generate reliable, effective XPaths for the web page.
Good case Bad case
Question Here’s a webpage on detail information with detail information of an NBA player. Please extract the height of the player. Here’s a webpage with detailed information about a university. Please extract the contact phone number of the university.
Case //div[@class=’gray200B-dyContent’]/ b[contains(text(),’Height:’)]/following-sibling::text() //div[@class=’infopage’]//h5[contains (text(), ’703-528-7809’)]
Table 4: Examples of XPath fragility. The green focuses on the common information across different webpages, while the red focuses on specific information of seed webpages.
Models Contains Equal(=)
GPT4 0.61% 2.90%
GPT-3.5-Turbo 9.33% 9.78%
Gemini Pro 10.62% 14.29%
Mixtral 8$\times$7B 12.88% 8.55%
Deepseek-Coder 11.63% 7.55%
CodeLlama 18.75% 14.29%
Mistral 7B 18.18% 33.33%
Table 5: Bad case ratio in two types of predicate.
XPath fragility within AutoCrawler
The fragility of XPath often refers to the characteristic of XPath expressions becoming ineffective or inaccurately matching the target element when faced with new webpages. This is mainly due to XPath specifying specific information through predicates, such as text, @class, etc.
We mainly focus on the fragility of text because these webpages are from the same websites (i.e. @class is a good characteristic for generating stable action sequences). Table 4 shows XPath expressions that rely on text. We aim to explore the reusability of generating XPath based on text features. We manually calculated the proportion of bad cases with two types of predicates, contains and equal  [^5] https://www.w3schools.com/xml/xpath_syntax.asp. The results in Table 5 show that the stronger LLMs capability, the lower the proportion of bad cases with AutoCrawler . However, it should be noted that the current SoTA LLM GPT-4 still suffers from an XPath fragility problem, which indicates that relying entirely on LLMs to generate reliable XPath still has some distance to go.
4.5 Error Analysis
We perform an analysis by looking at the recorded action sequence of AutoCrawler with GPT-4 and identify the following common failure modes. We mainly focus on the case that is categorized as unexecutable, over-estimate, and else.
Non-generalizability of webpages
The target information and corresponding webpage structures exhibit variations across different webpages, leading to a lack of generalizability in AutoCrawler (i.e., the inability to apply the same rules across all webpages in the same website) For instance, for the task "Please extract the name of the company that offers the job" in the website job-careerbuilder, most webpages contain the company name, but there is one webpage where the company name is listed as "Not Available" on another node of DOM tree.
Miss in multi-valued
Interesting Presented with the task of generating a crawler for extracting address in restaurant webpages or contact phone number from university websites, the target information is located in multiple locations in the webpage, such as the information bar, title, etc. Although AutoCrawler is capable of generating action sequences to extract portions of information, crafting a comprehensive action sequence that captures all of the information remains a challenge.
5 Related Work
5.1 Web Automation with LLMs
Many studies explore the concept of an open-world in web simulation environments Shi et al. (2017); Yao et al. (2023); Deng et al. (2023); Zhou et al. (2023), encompassing a broad spectrum of tasks found in real-life scenarios, such as online shopping, flight booking, and software development. Current web automation frameworks mainly aim to streamline the web environment Sridhar et al. (2023); Gur et al. (2023); Zheng et al. (2024) and to devise strategies for planning and interacting with the web Sodhi et al. (2023); Ma et al. (2023). However, these frameworks exhibit a lack of reusability, with agents heavily reliant on LLMs for even similar tasks, resulting in inefficiencies.
5.2 DOM-based Web Extraction
These methods utilize the hierarchical structure of the webpage. Method of this category includes rule-based Zheng et al. (2008), learning wrappers(i.e a DOM-specific parser that can extract content) Gulhane et al. (2011); Kushmerick (1997); Dalvi et al. (2011). Contemporary strategies employ distant supervision to autonomously create training samples by matching data from existing knowledge bases (KBs) with web sources Lockard et al. (2018, 2019). While this significantly lowers the effort required for annotation, it unavoidably leads to false negatives because of the incompleteness of knowledge bases (KBs) Xie et al. (2021).
6 Conclusion
In this paper, we introduce the crawler generation task and the paradigm that combines LLMs and crawlers to improve the reusability of the current web automation framework. We then propose AutoCrawler , a two-phase progressive understanding framework to generate a more stable and executable action sequence. Our comprehensive experiments demonstrate that AutoCrawler can outperform the state-of-the-art baseline in the crawler generation task.
Limitation
We introduce a paradigm that combines LLMs with crawlers for web crawler generation tasks and propose AutoCrawler to generate an executable action sequence with progressively understanding the HTML documents. Though experimental results show the effectiveness of our framework, there are still some limits to our work.
First, our framework is restricted to the paradigm in the information extraction task for vertical webpages. LLMs with crawlers provide high efficiency in open-world web IE tasks, but can hardly transfer to existing web environments such as Mind2Web Deng et al. (2023), WebArena Zhou et al. (2023).
Second, our framework rely on the performance of backbone LLMs. Enhancing LLMs’ ability to understand HTML is a very valuable research question, including corpus collection and training strategy. We will conduct research on HTML understanding enchancement in future work.
Ethic statement
We hereby declare that all authors of this article are aware of and adhere to the provided ACL Code of Ethics and honor the code of conduct.
Use of Human Annotations
Human annotations are only utilized in the early stages of methodological research to assess the feasibility of the proposed solution. All annotators have provided consent for the use of their data for research purposes. We guarantee the security of all annotators throughout the annotation process, and they are justly remunerated according to local standards. Human annotations are not employed during the evaluation of our method.
Risks
The datasets used in the paper have been obtained from public sources and anonymized to protect against any offensive information. Though we have taken measures to do so, we cannot guarantee that the datasets do not contain any socially harmful or toxic language.
Appendix A Experiments
A.1 Main results on Extended Swde
Because Extended Swde dataset focuses on OpenIE task (the relation is also expected to be extracted), we first map relations into a predefined list of attributes and remove unusual ones. Specifically, we conducted experiments with 294 attributes from 21 websites selected from the Extended Swde dataset.
Table D.2 shows the result. By comparing Table 3.1, we find that: 1) Under complex extraction task settings (multiple target values and ambiguous problem description), the closed-source LLMs perform better in generating executable action sequences compared to the open-source LLMs. 2) There are some tasks with unclear descriptions, such as the "Calendar System" and "Facilities and Programs Offered" on university websites, which affect the wrapper generation performance of all methods.
A.2 Main results on Ds1
Due to Ds1 only contains 166 hand-crafted webpages, and for each website, there are only two webpages, so we take one webpage for inference and the other for evaluation. Meanwhile, due to the number of the seed websites being equal to one, we test three methods without applying the synthesis module described in Section 3.3.
Table D.2 shows the result in the DS1 dataset. Among all LLMs with three methods, GPT4 + AutoCrawler achieves the best performance, and AutoCrawler beats the other two methods in all LLMs, which is consistent with the conclusion we make above.
Models Method Exec Eval IE Eval
Correct($\uparrow$) Unex.($\downarrow$) F1
GPT-3.5-Turbo COT 36.75 43.46 47.99
    - synthesis 27.56 57.24 34.44
[…]
    - synthesis 28.62 59.01 35.01
[…]
    - synthesis 44.52 29.33 58.44
Gemini Pro COT 29.69 47.19 41.81
    - synthesis 27.56 57.24 33.09
[…]
    - synthesis 28.62 59.01 37.60
[…]
    - synthesis 39.46 31.56 56.48
GPT4 COT 61.88 14.37 76.95
    - synthesis 46.88 30.00 61.20
[…]
    - synthesis 56.87 25.31 69.78
[…]
    - synthesis 65.31 11.87 80.41
Table 6: Ablation study on our two-stage framework. We report Correct, Unexecutable from the executive evaluation, and F1 score from the IE evaluation in Swde dataset.
Appendix B Analysis on AutoCrawler
B.1 Ablation Study
To further justify the effectiveness of each component of AutoCrawler , we perform an ablation study. The results are shown in Table A.2. It shows that: 1) AutoCrawlerwithout a second module still beat the other two baseline methods among different LLMs. 2) The second module of AutoCrawler , synthesis module, not only improves AutoCrawler , but also improves the performance of other methods. Using more webpages for inference can make the generated wrapper more stable and have better generalization.
Figure 4: The performance of GPT4 + AutoCrawler with different number of seed websites in Swde dataset.
B.2 Seed Websites
In all previous experiments, we fixed the number of seed websites $n_{s}=3$, which demonstrates the effectiveness of the synthesis module. In this experiment, we offer different numbers of seed webpages and test the performance of AutoCrawler with GPT4. The result is shown in Figure 4.
As the number of seed webpages increases, the correct ratio increases, while the unexecutable ratio decreases. It suggests that the performance of AutoCrawler can still be further improved by providing more seed webpages. In addition, the performance improvement reduces as the number increases, which shows that there is an upper limit to improve the performance of AutoCrawler by increasing the number of seed webpages.
B.3 Extraction Efficiency
Suppose the number of seed webpages is $n_{s}$, the number of webpages on the same website is $N_{\mathcal{W}}$, the time to generate a wrapper is $T_{g}$, the time to synthesis is $T_{s}$, and the time for extract information from a webpage with a wrapper is $T_{e}$. The total time for extracting all information from all websites with AutoCrawler is
$T_{1}=T_{G}+T_{E}=(n_{s}T_{g}+T_{s})+N_{\mathcal{W}}T_{e}$ (4)
Besides, the time for LLMs directly extracting information from a webpage is $T_{d}$, and the total time for extracting all information from all websites directly is
$T_{2}=N_{\mathcal{W}}T_{d}$ (5)
In real-world scenario, there are many webpages from the same websites to be extracted. Although generating a wrapper takes more time than extracting directly from a single webpage, the extraction efficiency of subsequent webpages would be significantly improved. To explore how many webpages are needed to make AutoCrawler more efficient in web IE, we calculate the threshold of $N_{\mathcal{W}}$. Suppose $T_{1}\leq T_{2}$, we have
$T_{G}+T_{E}=(n_{s}T_{g}+T_{s})+N_{\mathcal{W}}T_{e}\leq N_{\mathcal{W}}T_{d}$ (6)
$N_{\mathcal{W}}\geq\frac{n_{s}T_{g}+T_{s}}{T_{d}-T_{e}}$ (7)
It should be noted that $T_{g}$ depends on $d_{max}$ in Algorithm 1 and can be roughly considered as $T_{g}\approx d_{max}T_{d}$. In our experimental settings, we set $d_{max}=5$ and $n_{s}=3$. Also, under the approximation that $T_{s}\approx T_{d}$ and $T_{d}\gg T_{e}$, AutoCrawler have better extraction efficiency when a website contains more than 16 webpages.
Model F1
Render-Full Hao et al. (2011) 84.30
FreeDOM Lin et al. (2020) 82.32
SimpDOM Zhou et al. (2021) 83.06
[…]
WebFormer Wang et al. (2022) 92.46
Reflexion + GPT4 82.40
AutoCrawler + GPT4 88.69
Table 7: Comparing the extraction performance (Precision, Recall and F1) of 5 baseline models to our method AutoCrawler using GPT4 on the Swde dataset. Each value of the supervised model in the table is trained on 1 seed site.
B.4 Comparison with supervised baselines
Table 7 shows the comparison with 5 baseline models in web information extraction on supervised learning scenarios. These models are trained on webpages in some seed websites and tested on the other websites. Although the comparison is unfair because our methods is in zero-shot settings, AutoCrawler beat most of them on F1 scores. It shows that by designing an appropriate framework, LLMs can surpass supervised learning methods in some web information extraction tasks.
Figure 5: Comparison with the other two baselines.
B.5 Comparison with COT & Reflexion
Figure 5 more intuitively shows the specific differences between different baselines in the experiment. The most significant difference between AutoCrawler and other methods lies in whether the hierarchical structure of web pages is utilized to help LLMs reduce the difficulty of complex web structures. COT only executes one turn while the other executes multiple turns and can learn from the failed execution of the wrapper. Compared to the Reflexion method, AutoCrawler employs top-down and step-back operations to prune the DOM tree during each XPath generation process, thereby reducing the length of the web page. In contrast, the Reflextion method can only reflect and regenerate after producing an unexecutable XPath, which does not effectively simplify the webpage.
Appendix C Dataset Statistic
Table 101112 shows the detailed statistic about the semi-structure web information extraction dataset Swde, Extended Swde and Ds1.
Appendix D Prompt List
D.1 Task Prompt
Table 13 shows the task prompt we design for each attribute for Swde.
D.2 Module Prompt
We provide a comprehensive list of all the prompts that have been used in this study, offering a clear reference to understand our experimental approach.
Models Method Executable Evaluation IE Evaluation
Correct($\uparrow$) Prec Reca Unex.($\downarrow$) Over. Else Prec Reca F1
Closed-source LLMs
GPT-3.5-Turbo COT 34.49 3.48 4.53 56.10 0.35 1.05 87.96 42.16 40.58
Reflexion 43.90 1.74 2.09 49.13 0.35 2.79 93.46 49.58 48.66
AutoCrawler 45.30 4.18 8.01 35.89 0.35 6.27 83.60 60.84 56.69
[…]
Reflexion 34.15 2.09 6.97 51.57 0.35 4.88 84.43 45.19 41.66
AutoCrawler 35.89 5.23 10.10 42.86 0.35 5.57 83.74 52.75 47.73
[…]
Reflexion 63.76 3.83 5.57 20.91 0.35 5.57 86.00 76.50 74.50
AutoCrawler 63.07 3.48 5.92 16.72 0.35 10.45 81.29 78.77 74.77
Open-source LLMs
CodeLlama COT 9.01 1.29 2.15 85.84 0.00 1.72 87.22 12.62 11.21
Reflexion 13.73 1.72 3.00 80.26 0.00 1.29 89.41 17.76 16.01
AutoCrawler 11.16 0.00 1.72 85.84 0.00 1.29 92.49 13.29 12.52
[…]
Reflexion 29.62 1.05 4.18 62.02 0.35 2.79 83.44 36.44 33.64
AutoCrawler 40.07 3.83 9.41 39.37 0.35 6.97 81.63 57.10 51.57
[…]
Reflexion 36.24 3.48 3.83 51.92 0.00 4.53 83.53 45.03 43.64
AutoCrawler 37.63 2.44 5.92 50.52 0.35 3.14 86.91 47.09 44.33
Table 8: The executable evaluation and IE evaluation of LLMs with three frameworks in Extended Swde dataset. We examine 6 LLMs, including 3 closed-source LLMs and 3 open-source LLMs.
Models Method Executable Evaluation IE Evaluation
Correct($\uparrow$) Prec Reca Unex.($\downarrow$) Over. Else Prec Reca F1
Closed-source LLMs
GPT-3.5-Turbo COT 32.65 4.08 8.16 53.06 0.00 2.04 90.56 43.54 41.16
Reflexion 36.73 8.16 4.08 51.02 0.00 0.00 95.56 44.22 43.75
AutoCrawler 48.98 4.08 0.00 44.90 0.00 2.04 94.90 51.70 52.38
[…]
Reflexion 20.25 10.13 1.27 65.82 0.00 2.53 88.83 26.93 27.66
AutoCrawler 43.04 15.19 3.80 34.18 0.00 3.80 93.76 55.97 56.92
[…]
Reflexion 50.60 10.84 4.82 33.73 0.00 0.00 96.85 62.65 63.50
AutoCrawler 57.83 15.66 4.82 16.87 0.00 4.82 92.88 74.95 75.52
Open-source LLMs
CodeLlama COT 2.70 2.70 5.41 89.19 0.00 0.00 78.72 10.62 9.19
Reflexion 8.82 0.00 5.88 85.29 0.00 0.00 94.12 14.41 12.69
AutoCrawler 13.51 0.00 5.41 81.08 0.00 0.00 84.12 18.92 17.39
[…]
Reflexion 22.78 6.33 1.27 69.62 0.00 0.00 94.15 28.03 28.20
AutoCrawler 36.71 11.39 6.33 43.04 0.00 2.53 91.59 48.52 48.23
[…]
Reflexion 22.89 6.02 3.61 65.06 0.00 2.41 90.21 31.43 32.04
AutoCrawler 39.76 10.84 6.02 42.17 0.00 1.20 90.43 51.39 50.28
Table 9: The executable evaluation and IE evaluation of LLMs with three frameworks in Ds1 dataset. We examine 6 LLMs, including 3 closed-source LLMs and 3 open-source LLMs.
Listing 1: Prompts for ZS_COT, Reflexion and AutoCrawler
Prompt-crawler (for COT, Reflexion and AutoCrawler):
Please read the following HTML code, and then return an Xpath that can recognize the element in the HTML matching the instruction below.
Instruction: {0}
Here are some hints:
1. Do not output the xpath with exact value or element appears in the HTML.
2. Do not output the xpath that indicate multi node with different value. It would be appreciate to use more @class to identify different node that may share the same xpath expression.
3. If the HTML code doesnt contain the suitable information match the instruction, keep the xpath attrs blank.
4. Avoid using some string function such as substring()’ and normalize-space()’ to normalize the text in the node.
Please output in the following Json format:
"thought": "", # a brief thought of how to confirm the value and generate the xpath
"value": "", # the value extracted from the HTML that match the instruction
"xpath": "", # the xpath to extract the value
Heres the HTML code:
{1}
Prompt-Reflexion (for Reflexion and AutoCrawler):
Heres the HTML extraction task:
Task description: Please read the following HTML code, and then return an Xpath that can recognize the element in the HTML matching the instruction below.
Instruction: {0}
We will offer some history about the thought and the extraction result. Please reflect on the history trajectory and adjust the xpath rule for better and more exact extraction. Heres some hints:
1. Judge whether the results in the history is consistent with the expected value. Please pay attention for the following case:
1) Whether the extraction result contains some elements that is irrelevent
2) Whether the crawler return a empty result
3) The raw values containing redundant separators is considered as consistent because we will postprocess it.
2. Re-thinking the expected value and how to find it depend on xpath code
3. Generate a new or keep the origin xpath depend on the judgement and thinking following the hints:
1. Do not output the xpath with exact value or element appears in the HTML.
2. Do not output the xpath that indicate multi node with different value. It would be appreciate to use more @class and [num] to identify different node that may share the same xpath expression.
3. If the HTML code doesnt contain the suitable information match the instruction, keep the xpath attrs blank.
Please output in the following json format:
"thought": "", # thought of why the xpaths in history are not work and how to adjust the xpath
"consistent": "", # whether the extracted result is consistent with the expected value, return yes/no directly
"value": "", # the value extracted from the HTML that match the task description
"xpath": "", # a new xpath that is different from the xpath in the following history if not consistent
And heres the history about the thought, xpath and result extracted by crawler.
{1}
Heres the HTML code:
{2}
Prompt-synthesis (for COT, Reflexion and AutoCrawler):
Youre a perfect discriminator which is good at HTML understanding as well. Following the instruction, there are some action sequence written from several HTML and the corresponding result extracted from several HTML. Please choose one that can be best potentially adapted to the same extraction task on other webpage in the same websites. Here are the instruction of the task:
Instructions: {0}
The action sequences and the corresponding extracted results with different sequence on different webpage are as follow:
{1}
Please output the best action sequence in the following Json format:
"thought": "" # brief thinking about which to choose
"number": "" # the best action sequence choosen from the candidates, starts from 0. If there is none, output 0.
Prompt-judgement (for AutoCrawler):
Your main task is to judge whether the extracted value is consistent with the expected value, which is recognized beforehand. Please pay attention for the following case:
1) If the extracted result contains some elements that is not in expected value, or contains empty value, it is not consistent.
2) The raw values containing redundant separators is considered as consistent because we can postprocess it.
The extracted value is: {0}
The expected value is: {1}
Please output your judgement in the following Json format:
"thought": "", # a brief thinking about whether the extracted value is consistent with the expected value
"judgement": "" # return yes/no directly
Prompt-stepback (for AutoCrawler):
Your main task is to judge whether the following HTML code contains all the expected value, which is recognized beforehand.
Instruction: {0}
And heres the value: {1}
The HTML code is as follow:
{2}
Please output your judgement in the following Json format:
"thought": "", # a brief thinking about whether the HTML code contains expected value
"judgement": "" # whether the HTML code contains all extracted value. Return yes/no directly.
Domain Attribute Website Num Domain Attribute Website Num
Auto model price engine fuel_economy aol 2000 Movie ttitle director genre mpaa_rating allmovie 2000
autobytel 2000 amctv 2000
automotive 1999 boxofficemojo 2000
autoweb 2000 hollywood 2000
carquotes 2000 iheartmovies 2000
cars 657 imdb 2000
kbb 2000 metacritic 2000
motortrend 1267 msn 2000
msn 2000 rottentomatoes 2000
yahoo 2000 yahoo 2000
Book title author isbn_13 publisher pub_date abebooks 2000 NBAPlayer name team height weight espn 434
amazon 2000 fanhouse 446
barnesandnoble 2000 foxsports 425
bookdepository 2000 msnca 434
booksamillion 2000 nba 434
bookorders 2000 si 515
buy 2000 slam 423
christianbook 2000 usatoday 436
deepdiscount 2000 wiki 420
waterstone 2000 yahoo 438
Camera model price manufacturer amazon 1767 Restaurant name address phone cuisine fodors 2000
beachaudio 247 frommers 2000
buy 500 zagat 2000
compsource 430 gayot 2000
ecost 923 opentable 2000
jr 367 pickaretaurant 2000
newegg 220 restaurantica 2000
onsale 261 tripadvisor 2000
pcnation 234 urbanspoon 2000
thenerd 309 usdiners 2000
Job title company location date_posted careerbuilder 2000 University name phone website type collegeboard 2000
dice 2000 collegenavigator 2000
hotjobs 2000 collegeprowler 2000
job 2000 collegetoolkit 2000
jobcircle 2000 ecampustours 1063
jobtarget 2000 embark 2000
monster 2000 matchcollege 2000
nettemps 2000 princetonreview 615
rightitjobs 2000 studentaid 2000
techcentric 2000 usnews 1027
Table 10: Detail statistic of Swde dataset.
Domain Website # Attributes
Movie allmovie 20
amctv 13
hollywood 12
iheartmovies 8
imdb 34
metacritic 17
rottentomatoes 10
yahoo 10
NBAPlayer espn 10
fanhouse 14
foxsports 10
msnca 12
si 12
slam 12
usatoday 5
yahoo 9
University collegeprowler 18
ecampustours 14
embark 23
matchcollege 15
usnews 19
Table 11: Detail statistic of Extend Swde dataset.
Domain Attribute Website
Book title author price abebooks
alibris
barnesandnoble
fishpond
infibeam
powells
thriftbooks
E-commerce title price amazoncouk
bestbuy
dabs
ebay
pcworld
tesco
uttings
Hotel address price title agoda
expedia
hotels
hoteltravel
javago
kayak
ratestogo
venere
Movie actor genre title 123movieto
hollywoodreporter
imdb
mediastinger
metacritic
rottentomatoes
themoviedb
yidio
Table 12: Detail statistic of Ds1 dataset.
Domain Task prompt Prompt
Auto Here’s a webpage with detailed information about an auto. Please extract the model of the auto.
Please extract the price of the auto.
Please extract the engine of the auto.
Please extract the fuel efficiency of the auto.
Book Here’s a webpage with detailed information about a book. Please extract the title of the book.
Please extract the author of the book.
Please extract the isbn number of the book.
Please extract the publisher of the book.
Please extract the publication date of the book.
Camera Here’s a webpage with detail information of camera. Please extract the product name of the camera.
Please extract the sale price of the camera.
Please extract the manufacturer of the camera.
Job Here’s a webpage with detailed information about a job. Please extract the title of the job.
Please extract the name of the company that offers the job.
Please extract the working location of the job.
Please extract the date that post the job.
Movie Here’s a webpage with detailed information about a movie. Please extract the title of the movie.
Please extract the director of the movie.
Please extract the genre of the movie.
Please extract the MPAA rating of the movie.
NBAPlayer Here’s a webpage with detailed information about an NBA player. Please extract the name of the player.
Please extract the team of the player he plays now.
Please extract the height of the player.
Please extract the weight of the player.
Restaurant Here’s a webpage with detailed information about a restaurant. Please extract the restaurant’s name.
Please extract the restaurant’s address.
Please extract the restaurant’s phone number.
Please extract the cuisine that the restaurant offers.
University Here’s a webpage on detailed information about a university. Please extract the name of the university.
Please extract the contact phone number of the university.
Please extract the website url of the university.
Please extract the type of the university.
Table 13: Prompts for crawler generation task in Swde dataset.